The Motor Vehicle Department has found that the probability of a person passing the test for a drivers license on the first try is 0.74. The probability that an individual who fails on the first test will pass on the second try is 0.83, and the probability that an individual who fails the first and second tests will pass the third time is 0.65. Find the probabilities that an individual will do the following a. P(Fail both the first and second tests) b. P(Fail three times in a row) c. P(Require at least two tries) a. P(fails both the first and second tests) Type an integer or decimal rounded to four decimal places as needed.)
Answer #1

a)P(fail both first and second test) =(1-0.74)*(1-0.83)=0.0442

b)P(fail three times in a row) =(1-0.74)*(1-0.83)*(1-0.65)=0.01547

c)P(require at least two tries) =1-P(pass on first) =1-0.74 =0.26